# Jaime Angulo Pava  and  Alexander Muñoz

### v3

@@ 4.2. Boundary system for arbitrary amount of half-lines
For this second part we no longer seek for unitary dynamics. We feel confortable allowing the dynamics to be generated by a $C_{0}$-semigroup of contractions. To do so, we consider $\mathcal{G}$ consisting of finite amount $N$ of half-lines attached to the vertex $L$ with no restrictions on the signs of the coefficients of the Hamiltonian. We still treat $-L$ as a terminal vertex but we will be emphatic in whether the examples to be presented maintain or not the looping-edge structure. It is worth mentioning that most of the following results also apply to tadpole-like graphs.
@@
Consider the spaces $\mathcal{C}_{1}\cong\mathcal{C}_{0}\cong\mathbb{C}^{3(N+1)}$ and endow these spaces with the non-degenerated indefinite inner products
$\left(x\mid y\right)_{1}:=\left(C_{1}x\mid y\right)\mbox{ and }\ \left(x\mid y\right)_{0}:=\left(C_{0}x\mid y\right),$ (4.24)
@@
$C_{1}:=\begin{pmatrix}B_{0}\\ &B_{1}\\ &&B_{2}\\ &&&\ddots\\ &&&&B_{N}\end{pmatrix}\mbox{ and }C_{0}:=\frac{1}{N+1}\begin{pmatrix}B_{0}\\ &B_{0}\\ &&B_{0}\\ &&&\ddots\\ &&&&B_{0}\end{pmatrix}.$
$(\mathcal{C}_{1},\left(\cdot\mid\cdot\right)_{1})$ and $(\mathcal{C}_{0},\left(\cdot\mid\cdot\right)_{0})$ are Krein spaces.
Under these new notations, Proposition 4.1 can be rewritten as
$[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}-\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}.$ (4.25)
@@
We consider the function
$\begin{split}\mathcal{F}:&\ G(A_{0}^{*})\to\mathcal{C}_{1}\oplus\mathcal{C}_{0}\\ &(\textbf{U},A_{0}^{*}\textbf{U})\mapsto\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix},\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)\end{split}$ (4.26)
4.2.1. Extensions generating contractile dynamics
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be a densely defined operator. Consider an extension $A_{\mathcal{L}}$ of $A_{0}$ defined by the relation $G(A_{\mathcal{L}})=\mathcal{F}^{-1}(G(\mathcal{L}))$. In other words,
$D(A_{\mathcal{L}}):=\left\{\textbf{U}\in D(A_{0}^{*})\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\in D(\mathcal{L})\mbox{ and }\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right\}.$ (4.27)
acting as $A_{\mathcal{L}}\textbf{U}:=-A_{0}^{*}\textbf{U}$.
Let us determine the adjoint of operator of $A_{\mathcal{L}}$.
@@ 4.2.1. Extensions generating contractile dynamics
For $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ densely defined we have
$D(A_{\mathcal{L}}^{*}):=\left\{(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{0}^{*})\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\in D(\mathcal{L}^{\#})\mbox{ and }\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right\}$ (4.28)
acting as $A_{\mathcal{L}}^{*}\textbf{U}=A_{0}^{*}\textbf{U}.$
@@
Let $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{\mathcal{L}})$, $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})\in D(A_{0}^{*})$. From (4.7) and the definition of $A_{\mathcal{L}}$ we have
$\begin{split}[A_{\mathcal{L}}\textbf{U},\textbf{V}]&=[\textbf{U},A_{0}^{*}\textbf{V}]+\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}\\ &=[\textbf{U},A_{0}^{*}\textbf{V}]+\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}.\end{split}$ (4.29)
Therefore, $\textbf{V}\in D(A_{\mathcal{L}}^{*})$ if and only if for all $\textbf{U}\in D(A_{\mathcal{L}})$ it follows
$\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}.$ (4.30)
Note from the definition of $A_{\mathcal{L}}$ that $\mathcal{F}(G(A_{\mathcal{L}}))=G(\mathcal{L})$; which means
$D(\mathcal{L})=\left\{\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\textbf{U}\in D(A_{\mathcal{L}})\right\}.$ (4.31)
We conclude (4.30) is equivalent to
#$\mathcal{L}^{\#}\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}=\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}.$
From Proposition 4.4 we see that knowing the particular expression defining $\mathcal{L}^{\#}$ is important to determine explicitly the domain of $A_{\mathcal{L}}^{*}$. Note in case $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ is given by a $3(N+1)\times 3(N+1)$ matrix, the $(\mathcal{C}_{0},\mathcal{C}_{1})$-adjoint $\mathcal{L}^{\#}$ is given by the expression
#$\mathcal{L}^{\#}=C_{1}^{-1}\mathcal{L}^{*}C_{0}.$ (4.32)
In fact, for $y\in\mathcal{C}_{0}$ the definition of $\mathcal{L}^{\#}$ imply that for all $x\in\mathcal{C}_{1}$ we have
$\left(C_{0}\mathcal{L}x\mid y\right)=\left(\mathcal{L}x\mid y\right)_{0}=\left(x\mid\mathcal{L}^{\#}y\right)_{1}=\left(C_{1}x\mid\mathcal{L}^{\#}y\right)_{0}.$ (4.33)
Since $C_{1}$ is real symmetric we obtain that for all $x\in\mathcal{C}_{1}$
$\left(x\mid\ \left(\mathcal{L}^{*}C_{0}-C_{1}\mathcal{L}^{\#}\right)y\right)=0.$ (4.34)
Since $y$ is arbitrary, it follows
$\mathcal{L}^{*}C_{0}=C_{1}\mathcal{L}^{\#},$ (4.35)
which implies (4.32)
In the following two propositions we characterize when both, $A_{\mathcal{L}}$ and $A_{\mathcal{L}}^{*}$, are dissipative operators. We recall that dissipative operators on Hilbert spaces were briefly described on subsection 2.2.
@@
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$. $A_{\mathcal{L}}$ is dissipative if and only if $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction.

### v4

@@ 4.2. Boundary system for arbitrary amount of half-lines
For this part we no longer seek unitary dynamics. We allow the dynamics to be generated by a $C_{0}$-semigroup of contractions. We consider $\mathcal{G}$ consisting of a finite amount $N$ of half-lines attached to the vertex $L$, with no restrictions on the signs of the coefficients of the Hamiltonian. We still virtually treat L-L as a terminal vertex.
@@
Consider the spaces $\mathcal{C}_{1}\cong\mathcal{C}_{0}\cong\mathbb{C}^{3(N+1)}$ equipped with the non-degenerate indefinite inner products
$(x\mid y)_{1}:=(C_{1}\,x\mid y),\qquad(x\mid y)_{0}:=(C_{0}\,x\mid y),$ (4.24)
@@
$C_{1}:=\operatorname{diag}(B_{0},B_{1},B_{2},\dots,B_{N}),\qquad C_{0}:=\frac{1}{N+1}\operatorname{diag}(B_{0},B_{0},\dots,B_{0}).$
Both $(\mathcal{C}_{1},(\cdot\mid\cdot)_{1})$ and $(\mathcal{C}_{0},(\cdot\mid\cdot)_{0})$ are Krein spaces.
4.2.1. Boundary trace maps and the Green identity
We introduce boundary trace maps $\Gamma_{1},\Gamma_{0}:D(A_{0}^{*})\to\mathbb{C}^{3(N+1)}$ by
$\Gamma_{1}\mathbf{U}:=\begin{pmatrix}\partial\phi(-L)\\[2.0pt] \partial\psi_{1}(L)\\ \vdots\\ \partial\psi_{N}(L)\end{pmatrix}\in\mathcal{C}_{1},\qquad\Gamma_{0}\mathbf{U}:=\begin{pmatrix}\partial\phi(L)\\ \partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\in\mathcal{C}_{0}.$ (4.25)
In terms of these maps, the Green identity for $A_{0}^{*}$ reads
$A_{0}^{*}\mathbf{U},\mathbf{V}]+[\mathbf{U},A_{0}^{*}\mathbf{V}]=(\Gamma_{1}\mathbf{U}\mid\Gamma_{1}\mathbf{V})_{1}-(\Gamma_{0}\mathbf{U}\mid\Gamma_{0}\mathbf{V})_{0}.$ (4.26)
4.2.2. Boundary relations, Krein adjoints, and contractive extensions
@@
A boundary relation is a linear subspace $R\subset\mathcal{C}_{1}\times\mathcal{C}_{0}$. Its Krein adjoint $R^{[*]}\subset\mathcal{C}_{0}\times\mathcal{C}_{1}$ is defined by
$(y,x)\in R^{[*]}\;\Longleftrightarrow\;(y_{1}\mid y)_{0}=(x_{1}\mid x)_{1}\quad\text{for all }(x_{1},y_{1})\in R.$ (4.27)
The relation $R$ is called a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction if
[…]
When $R=\operatorname{graph}(\mathcal{L})$ for a linear map $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$, the Krein adjoint satisfies $R^{[*]}=\operatorname{graph}(\mathcal{L}^{\#})$ with
$\mathcal{L}^{\#}:=C_{1}^{-1}\mathcal{L}^{*}C_{0},$ (4.29)
and contractivity of $R$ reduces to $\mathcal{L}^{*}C_{0}\mathcal{L}-C_{1}\preceq 0$.
Remark 1.
The multivalued part of $R$ is $\operatorname{mul}(R):=\{y:(0,y)\in R\}$. Whenever $\operatorname{mul}(R)\neq\{0\}$, the relation $R$ is not the graph of any operator, and no matrix $\mathcal{L}$ can encode the corresponding boundary conditions. Concretely, $(0,y)\in R$ imposes $\Gamma_{1}\mathbf{U}=0$, i.e.
$\phi(-L)=\phi^{\prime}(-L)=\phi^{\prime\prime}(-L)=0,\quad\psi_{j}(L)=\psi_{j}^{\prime}(L)=\psi_{j}^{\prime\prime}(L)=0,\quad j=1,\dots,N,$
while leaving $\Gamma_{0}\mathbf{U}$ unconstrained. These are Dirichlet-type conditions at the incoming edges; they are valid generators of contraction semigroups whenever the corresponding relation is contractive, but are inaccessible to the operator-based framework.
Definition 4.6.
Given a boundary relation $R\subset\mathcal{C}_{1}\times\mathcal{C}_{0}$, define
$A_{R}\mathbf{U}:=-A_{0}^{*}\mathbf{U},\qquad D(A_{R}):=\bigl\{\mathbf{U}\in D(A_{0}^{*}):(\Gamma_{1}\mathbf{U},\Gamma_{0}\mathbf{U})\in R\bigr\}.$ (4.30)
Let us describe the adjoint of $A_{R}$.
@@
$A_{R}^{*}=-A_{R^{[*]}}$, where $A_{R^{[*]}}$ acts as $A_{0}^{*}$ with domain
$D\!\left(A_{R^{[*]}}\right):=\bigl\{\mathbf{V}\in D(A_{0}^{*}):(\Gamma_{0}\mathbf{V},\Gamma_{1}\mathbf{V})\in R^{[*]}\bigr\}.$ (4.31)
@@
Let $\mathbf{U}\in D(A_{R})$ and $\mathbf{V}\in D(A_{0}^{*})$. Using (4.26) and $A_{R}\mathbf{U}=-A_{0}^{*}\mathbf{U}$,
$A_{R}\mathbf{U},\mathbf{V}]=[\mathbf{U},A_{0}^{*}\mathbf{V}]+(\Gamma_{0}\mathbf{U}\mid\Gamma_{0}\mathbf{V})_{0}-(\Gamma_{1}\mathbf{U}\mid\Gamma_{1}\mathbf{V})_{1}.$ (4.32)
Hence $\mathbf{V}\in D(A_{R}^{*})$ if and only if
$(y_{1}\mid\Gamma_{0}\mathbf{V})_{0}=(x_{1}\mid\Gamma_{1}\mathbf{V})_{1}\quad\forall\,(x_{1},y_{1})\in R,$
which by (4.27) is equivalent to $(\Gamma_{0}\mathbf{V},\Gamma_{1}\mathbf{V})\in R^{[*]}$. In this case $A_{R}^{*}\mathbf{V}=A_{0}^{*}\mathbf{V}=-(-A_{0}^{*}\mathbf{V})=-A_{R^{[*]}}\mathbf{V}$. ∎
We now obtain dissipativity via contractivity in the following sense:
@@
$A_{R}$ is dissipative if and only if $R$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction.